# ShanghaiTech Mapping Robot is All You Need: Robot System for Collecting Universal Ground Vehicle Datasets

### v2

@@ VI Calibration
To ensure accurate perception and mapping, it is crucial to calibrate the intrinsic and extrinsic parameters of the various sensors employed in the system. There are various tools suitable for our use case, including ones concentrate on sensors of one same type [42], [43], [44], and others performing cross-modal calibration [45], [46], [47] and [48]. This section outlines the calibration procedures for the RGB cameras, event-based cameras, and LiDAR sensors.
@@ VI-A Camera Intrinsic & Extrinsic Calibration
The intrinsic and extrinsic parameters of 21 out of the 22 RGB cameras, including 10 Grasshopper3, 5 Azure Kinect DK, and 6 individual sensors of the LadyBug5+, were calibrated using the multi-camera calibration toolbox Kalibr [42] and AprilGrid patterns. Kalibr is capable of calibrating the intrinsic and extrinsic parameters of camera systems with non-globally shared overlapping fields of view and supports various distortion models, such as radial-tangential and equidistant models, which accurately represent the lenses used on the Grasshopper3 and the built-in fisheye lenses on the Ladybug5+, respectively.
For each of the 5 main directions (front, back, left, right, and up), the 2 Grasshopper3s, 1 Azure Kinect DK, and 1 or 2 sensors of the LadyBug5+ facing that side are grouped and calibrated together since they share a similar perspective. This gave us intrinsics of all cameras and 5 extrinsics groups, each independent from the others. Subsequently, the extrinsics between the 5 surrounding LadyBug5+ sensors were calibrated in a separate run, connecting the extrinsics of the 4 horizontal sides (front, back, left, and right). The up-looking side was finally connected to the extrinsic tree using the factory pre-calibrated transformation between the front-looking and up-looking sensors of the LadyBug5+.
Figure 18: Reprojection error of 4 frontward looking RGB cameras after calibration by Kalibr. (cam0: stereo left Grasshopper3, cam1: stereo right Grasshopper3, cam2: Azure Kinect DK RGB sensor, cam3: LadyBug5+ sensor number 0)
The intrinsic parameters of the last, down-looking Grasshopper3 camera were calibrated using the ROS camera calibration package [49] and a checkerboard target with a size matching the small area of view on the ground. Considering that the camera has no overlapping field of view with any other onboard sensors and that the down-looking image stream is more likely to be used in monocular image registration algorithms rather than visual odometry algorithms, its extrinsic parameters were derived from the structure design instead of using complex calibration algorithms.
@@ VI-C Calibration of Other Sensors
The MTi-630R IMU’s Allan Variance parameters, including random walk and bias instability, were calibrated using the allan_variance_ros tool [51]. With the IMU intrinsics, the Kalibr toolbox calibrated the extrinsic between the IMU and the 2 front-looking Grasshopper3 cameras. The mmWave radars on each side share an identical extrinsic with the respective Azure Kinect camera they attached onto, derived from the 3D printed mount’s design. The extrinsics of the ultrasound range finders and other unmentioned sensors were either roughly measured manually or derived from their mounts’ structure design.
Figure 19: Coordinates of calibrated sensors. Purple lines indicate the edges of transformation tree. An image of mapping robot took from a roughly same perspective is also shown.
@@ VII Procedures and Post-Processing
The mapping robot is designed with distributed computation in mind so it is vital to orchestrate all the nodes in the computation cluster along with the sensors connected to them. In this section we would introduce the procedure of data acquisition and post-process to be performed on the collected data.
@@ VII-A1 Material Preparation
At minimum, the robot itself, one battery, and a controlling terminal like KVM or laptop is needed for performing a acquisition run. A second battery as well as backup ones or power adaptor is recommended.
@@ VII-A2 Power Up
Close the main breaker on the power box. Husky UGV, the master node of the cluster, Syncboard as well as sensors getting electricity from the power box mentioned in Sec. IV-D will be powered up. Following the boot of master node, the robot becomes maneuverable.
@@ VII-A3 Cluster Bringup
The master node sets itself up as the Ethernet gateway and network file systems for Boot-from-LAN, then wakes up all the worker nodes. After they boot, we can control the nodes and access data stored in them. Network access also becomes available for nodes, Syncboard and GNSS receiver.
@@ VII-A4 Time Synchronization
Syncboard gets time base from external master clock and broadcasts it over cluster LAN. We would now perform a manual verification of nodes’ clock, then launch the triggering for sensor test.
@@ VII-A5 Sensors Bringup
The master node orchestrates all worker nodes to launch the sensor drivers. We would manually check if all sensors work as intended. If that’s not the case, sensor’s parameter would be modified accordingly. After everything is confirmed to work well, the triggering is turned off await the upcoming data acquisition.
@@ VII-A6 Data acquisition
Command the master node to orchestrate a distributed recording of sensor data. Turn on triggering to officially start a data acquisition run. We would now drive the robot along the data acquisition route.
@@ VII-A7 End of acquisition
After the robot reaches finishing position of the route, the triggering is turned off first, followed by the data recording processes. Sensor drivers stay running and do not have to be spawned repetitively until there is no more data acquisition runs to be carried out before powering off the robot. The data collected can be copied either into external hard drives via USB port from each nodes or to a workstation or network storage device through Ethernet connection, pending for post-processing.
@@ VII-B1 Timestamp Restoration
Replace the timestamps in sensor messages captured by hardware triggered sensors (Grasshopper3 camera, MTi-630R IMU, etc.) with precise triggering time of Syncboard. Any frame drop would be discovered and evaluated at the same time.
@@ VII-B2 Bayer Interleaving
LadyBug5+ on-the-fly compresses each image for the 6 camera sensors in 4 separate Bayer channels using JPEG, resulting in 24 JPEG images per one capture. We decompress and interleave them to get 6 image streams, one sensor each, in bgr8 format.
@@ VII-B3 Image Compression
To provide an alternative compressed version for faster distribution of our dataset, we encode each image stream using AV1 codec [52].
@@ VII-B4 Mapping Ground Truth
A FARO Focus 3D scanner is used to obtain a precise point cloud of the environment along the data acquisition route, acting as the mapping or 3D reconstruction ground truth.
@@ VII-B5 Trajectory Ground Truth
For route sections where GNSS signal stays available, we will post-process the onboard differential GNSS/INS data to obtain a ground truth trajectory with a typical precision of 1cm. Other than that, we register LiDAR scans to the mapping ground truth to get pose of robot as an alternative ground truth for indoor datasets.
VIII LiDAR&RGBD Interference Evaluation
LiDARs and RGB-Depth cameras on the mapping robot emit infra-red beams and measure the time of flight to obtain distance readings. There are concerns about potential interference between the sensors. In this experiment, we collected point clouds and depth images at a fixed position against a flat, white wall in the basement of ShanghaiTech University with each LiDAR and two RGB-Depth cameras working alone (sole runs) and together (collective run) for 10 seconds, as showed in Fig. 20.
Figure 20: Interference experiment scene setup
@@ VIII-A Precision analysis
TABLE II: Mean & S.D. of temporal S.D. of distance (m)
@@
Fig. 2121(a) & 21(b) shows the temporal mean of the point cloud collected by the front Robosense Bpearl. In the collective run, both Robosense Bpearl and Livox Avia give noisy points behind the wall or below the floor, indicating interference. However, the non-noisy points have consistent standard deviation of distance compared to the sole run, suggesting that with proper filtering, the two sensors would give point clouds of the same precision regardless of interference.
@@ VIII-B Accuracy analysis
TABLE III: Mean & S.D. of Point-to-Plane distance (m)
@@
TABLE IV: Pixel-wise Mean & S.D. of Distance Difference(m)
@@
For depth cameras, the difference in average distance measurements between the sole and collective runs is calculated pixel-wise to detect accuracy changes. As showed in Table IV, the mean difference is much smaller than its standard deviation, indicating that the depth cameras’ accuracy is also unaffected by interference.
@@ IX Outdoor Dataset
To assess the mapping robot’s data acquisition capability and verify the compatibility of the collected data with SLAM algorithms, we conducted an outdoor experiment on the ShanghaiTech University campus. The experiment involved collecting data along a predefined route and running SLAM algorithms on the collected data.
@@ IX-A Data Acquisition
An 1.4 km data acquisition route, as shown in Fig. 23, was designed to include a variety of terrains and scenes on the ShanghaiTech University campus. The route was evenly split between areas around buildings and parkland, with the mapping robot primarily driving on paved roads or solid structures like bridges, and briefly on grass. Three sections of the route, surrounded by buildings, woods, or a blend of both, were revisited after traveling elsewhere, providing loop-closing opportunities for SLAM algorithms. The route also included slopes to enrich the elevation diversity.
Figure 22: Mapping robot performing data acquisition on ShanghaiTech campus
The experiment was conducted on an overcast day, providing even lighting on objects and preventing direct sunlight spots in the collected images. The robot took 26 minutes to traverse the entire route, ultimately collecting 2,618 GiB of data, as shown in Fig. 23. A detailed list of the recorded ROS topics is provided in the Appendix A.
@@ IX-C Visual SLAM
X Conclusions
XI Acknowledgments
This work was supported by Science and Technology Commission of Shanghai Municipality (STCSM), project 22JC1410700 “Evaluation of real-time localization and mapping algorithms for intelligent robots”. This work has also been partially funded by the Shanghai Frontiers Science Center of Human-centered Artificial Intelligence.
Appendix A Details of ROS topics recorded in experiment
In Table. V, we showed for each type and model of sensor the message topics they publish to as well as the total number of messages recorded from them. The time duration, data rate, and message type are also listed.
TABLE V: Sensor Data Statistics of Outdoor Dataset

### v3

@@ VI Calibration
To ensure accurate perception and mapping, it is crucial to calibrate the intrinsic and extrinsic parameters of the various sensors employed in the system. There are various tools suitable for our use case, including ones that concentrate on sensors of one same type [44], [45], [46], and others performing cross-modal calibration [47], [48], [49], [50], [51] and [52]. This section outlines the calibration procedures for the RGB cameras, event-based cameras, and LiDAR sensors.
@@ VI-A Camera Intrinsic & Extrinsic Calibration
The intrinsic and extrinsic parameters of 21 out of the 22 RGB cameras, including 10 Grasshopper3, 5 Azure Kinect DK, and 6 individual sensors of the LadyBug5+, but excluding the downlooking Grasshopper3, were calibrated using the multi-camera calibration toolbox Kalibr [44] and AprilGrid patterns. Kalibr is capable of calibrating the intrinsic and extrinsic parameters of camera systems with non-globally shared overlapping fields of view and supports various distortion models, such as radial-tangential and equidistant models, which accurately represent the lenses used on the Grasshopper3 and the built-in fisheye lenses on the Ladybug5+.
For each of the 5 main directions (front, back, left, right, and up), the two Grasshopper3s, one Azure Kinect DK, and one or two sensors of the LadyBug5+ facing that side are grouped and calibrated together since they share a similar perspective. This gives us intrinsics of all cameras and 5 extrinsics groups, each independent from the others. Subsequently, the extrinsics between the 5 surrounding LadyBug5+ sensors were calibrated in a separate run, connecting the extrinsics of the 4 horizontal sides (front, back, left, and right). Because all up-looking and down-looking cameras do not share sufficient overlap with any side-looking cameras, Kalibr cannot be used to estimate their extrinsics. Thus the up-looking side was finally connected to the extrinsic tree using the factory pre-calibrated transformation between the front-looking and up-looking sensors of the LadyBug5+.
Figure 18: Reprojection error of 4 frontward looking RGB cameras after calibration by Kalibr. (cam0: stereo left Grasshopper3, cam1: stereo right Grasshopper3, cam2: Azure Kinect DK RGB sensor, cam3: LadyBug5+ sensor number 0).
The intrinsic parameters of the last, down-looking Grasshopper3 camera were calibrated using the ROS camera calibration package [53] and a checkerboard target with a size matching the small area of view on the ground. Considering that the camera has no overlapping field of view with any other onboard sensors and that it will not be used for 3D SLAM but rather down-looking odometry, its extrinsic parameters were derived from the structure design instead of using complex calibration algorithms.
@@ VI-C Calibration of Other Sensors
The MTi-630R IMU’s Allan Variance parameters, including random walk and bias instability, were calibrated using the allan_variance_ros tool [55]. With the IMU intrinsics, the Kalibr toolbox calibrated the extrinsic between the IMU and the two front-looking Grasshopper3 cameras. The mmWave radars on each side share an identical extrinsic with the respective Azure Kinect camera they are attached onto, derived from the 3D printed mount’s design. The extrinsics of the ultrasound range finders and other unmentioned sensors were either roughly measured manually or derived from their mounts’ structure design.
Figure 19: Coordinates of calibrated sensors. Purple lines indicate the edges of transformation tree. An image of mapping robot taken from a similar perspective is also shown.
@@ VII Procedures and Post-Processing
The mapping robot is designed with distributed computation in mind, so it is vital to orchestrate all the nodes in the computation cluster along with the sensors connected to them. In this section we introduce the procedure of data acquisition and post-processing to be performed on the collected data.
@@ VII-A1 Material Preparation
At minimum, the robot itself, one battery, and a controlling terminal like KVM (Keyboard VGA Mouse) or laptop is needed for performing a acquisition run. A second battery as well as backup ones or power adaptor is recommended.
@@ VII-A2 Power Up
Closing the main breaker on the power box, the Husky UGV, the nodes of the cluster, the Syncboard as well as sensors will be provided with electricity from the battery, as mentioned in Sec. IV-D. Only the master node (Node 1) and Syncboard boot when power becomes available. After Node 1 boot process is finished, the robot becomes maneuverable.
@@ VII-A3 Cluster Bringup
The master node sets itself up as the Ethernet gateway and network file systems for Boot-from-LAN, then wakes up all the worker nodes. Once they have booted up, we can control the nodes and access data stored in them. Network access also becomes available for nodes, the Syncboard and the GNSS receiver.
@@ VII-A4 Time Synchronization
The Syncboard gets the time base from an external master clock and broadcasts it over the cluster LAN. We use scripts to trigger immediate time synchronization on the nodes and print their clock times for manual verification, then launch the triggering for sensor testing procedure of the next step.
@@ VII-A5 Sensors Bringup
The master node orchestrates all worker nodes to launch the sensor drivers. We would manually check if all sensors work as intended. If that’s not the case, sensor’s parameters would be modified accordingly. After everything is confirmed to work well, the triggering is turned off awaiting the upcoming data acquisition.
@@ VII-A6 Data Acquisition
Here we command the master node to orchestrate a distributed recording of sensor data. Then we turn on triggering to officially start a data acquisition run. We would now drive the robot along the data acquisition route.
@@ VII-A7 End of Acquisition
After the robot reaches the finishing position of the route, the triggering is turned off first, followed by the termination of the data recording processes. Sensor drivers stay running and do not have to be spawned repetitively until there is no more data acquisition runs to be carried out before powering off the robot. The data collected can be copied either into external hard drives via a USB port from each node or to a workstation or network storage device through Ethernet connection, pending for post-processing.
@@ VII-B1 Timestamp Restoration
In order to be able to utilize the actual trigger times of the sensor data, which is important for many algorithms like stereo calculation or sensor fusion, especially with the IMUs working at a high frequency, we re-timestamp all those messages with the time of the actual trigger, as recorded by the Syncboard, rather than the time the data was processed on the node. This is done for the messages from all hardware-triggered sensors (Grasshopper3, LadyBug5+, Kinect Azure and MTi-630R IMU) shown in Fig. 17. Any frame drop would be discovered and noted at the same time.
@@ VII-B2 Bayer Interleaving
LadyBug5+ on-the-fly compresses each image for the 6 camera sensors in 4 separate Bayer channels using JPEG, resulting in 24 JPEG images per one capture. We decompress and interleave them to get 6 image streams, one for each sensor, in bayer_rggb8 format.
@@ VII-B3 Image Compression
To provide an alternative compressed version for faster distribution of our dataset, we debayer each image stream into YUV format and compress them with our ROS to AV1 encoding tool [56].
@@ VII-B4 Mapping Ground Truth
As an optional step, for acquiring 3D ground truth point clouds, especially also for estimating the indoor ground truth robot trajectory, a FARO Focus 3D scanner can be used to obtain a precise point cloud of the environment along the data acquisition route, acting as the mapping or 3D reconstruction ground truth.
@@ VII-B5 Trajectory Ground Truth
For route sections where GNSS signal is available, we will post-process the onboard differential GNSS/INS data to obtain a ground truth trajectory with a typical precision of 1cm. Other than that, we register LiDAR scans to the mapping ground truth to get the pose of robot as an alternative ground truth for indoor datasets.
VIII LiDAR & RGB-D Interference Evaluation
LiDARs and RGB-Depth cameras on the mapping robot emit infra-red beams and measure the time of flight to obtain distance readings. There are concerns about potential interference between the sensors. In this experiment, we collected point clouds and depth images at a fixed position against a flat, white wall in the basement of ShanghaiTech University with each LiDAR and two RGB-Depth cameras working alone (sole runs) and together (collective run) for 10 seconds, as shown in Fig. 20.
Figure 20: Interference experiment scene setup.
@@ VIII-A Precision Analysis
TABLE II: Interference Evaluation: Mean & S.D. of temporal S.D. of distance
@@
Fig. 2121(a) & 21(b) shows the temporal mean of the point cloud collected by the front Robosense Bpearl. In the collective run, both Robosense Bpearl and Livox Avia give noisy points behind the wall or below the floor, indicating interference. However, all valid objects are still detected by the LiDARs with substantial number of non-noisy points that have consistent standard deviation of distance compared to the sole run, suggesting that with proper filtering, the two sensors would give point clouds of the same precision regardless of interference.
@@ VIII-B Accuracy analysis
TABLE III: Interference Evaluation: Mean & S.D. of Point-to-Plane distance
@@
TABLE IV: Interference Evaluation:
Pixel-wise Mean & S.D. of Distance Difference
@@
For depth cameras, the difference in average distance measurements between the sole and collective runs is calculated pixel-wise to detect accuracy changes. As shown in Table IV, the mean difference is much smaller than its standard deviation, indicating that the depth cameras’ accuracy is also unaffected by interference.
@@ IX Outdoor Dataset
To assess the mapping robot’s data acquisition capability and verify the compatibility of the collected data with SLAM algorithms, we conducted an outdoor experiment on the ShanghaiTech University campus. The experiment involved collecting data along a predefined route and then running SLAM algorithms on the collected data.
@@ IX-A Data Acquisition
A 1.4 km data acquisition route, as shown in Fig. 23, was designed to include a variety of terrains and scenes on the ShanghaiTech University campus. The route was evenly split between areas around buildings and parkland, with the mapping robot primarily driving on paved roads or solid structures like bridges, and briefly on grass. Three sections of the route, surrounded by buildings, woods, or a blend of both, were revisited after traveling elsewhere, providing loop-closing opportunities for SLAM algorithms. The route also included slopes to enrich the elevation diversity.
Figure 22: Mapping robot performing data acquisition on ShanghaiTech campus.
The experiment was conducted on a quite overcast day, providing even lighting on objects and preventing direct sunlight spots in the collected images. The robot took 26 minutes to traverse the entire route, ultimately collecting 2,618 GiB of data, as shown in Fig. 23. A detailed list of the recorded ROS topics is provided in Appendix A.
@@ IX-B LiDAR-based SLAM
Our LiDAR SLAM implementation integrates two state-of-the-art algorithms: LIO-SAM [57] and MA-LIO [58]. For LIO-SAM, we activate loop closure detection and utilize the front-mounted RoboSense Ruby LiDAR sensor. Fig. 24 provides an example of the LiDAR point cloud map generated using LIO-SAM. In contrast, the MA-LIO configuration operates without loop closure, instead leveraging data from multiple LiDAR units, including RoboSense and Hesai sensors.
As demonstrated in Fig. 25, the multi-sensor fusion with the vertical Robosense Ruby LiDAR allows the system to accurately capture the facades of high-rise buildings and the sides of towers. The expanded sensor array enabled the reconstruction of a more complete 3D map, which proves especially beneficial in urban areas with prominent vertical structures.
Figure 24: LiDAR Point Cloud Map Generated Using LIO-SAM.
Figure 25: Mapping the ShanghaiTech Tower: MA-LIO without vertical LiDAR (left), MA-LIO with vertical LiDAR (middle), and photo (right).
@@ IX-C Visual SLAM
For visual SLAM, we employ ORB-SLAM3 [32] in both monocular-inertial and stereo-inertial configurations. Our setup utilizes the front-left camera for monocular mode and the front stereo camera pair for stereo mode. We conduct a comparative analysis of the results from these two configurations to evaluate their relative performance.
We use evo [59] to visualize and compare the pose estimation results from different SLAM algorithms. For visualization, we apply evo’s alignment function to adjust both the translation and scale of the estimated trajectories relative to the ground truth from the RTK system. Fig. 26 shows the aligned trajectory produced by LIO-SAM, MA-LIO, and ORB-SLAM3, overlaid with the ground truth trajectory from the RTK system. Fig. 27 provides a detailed breakdown of the trajectory in the x, y, and z dimensions, offering deeper insight into the variations in pose estimation over time across the different SLAM systems. Table V compares the absolute pose error (APE) and relative pose error (RPE) of LIO-SAM, MA-LIO, and two configurations of ORB-SLAM3, underscoring the superior performance of LiDAR-based methods in environments with significant visual challenges.
Despite the assistance of an inertial measurement unit (IMU), visual SLAM systems such as ORB-SLAM3 still suffer from scale drift, particularly in monocular configurations where depth information is lacking. While the stereo-inertial setup mitigates this to some extent by providing direct depth measurements, scale drift remains a challenge, especially in long-duration or visually-degraded scenarios. LiDAR-based SLAM methods, by contrast, exhibit far more stable performance in such environments due to the precise distance measurements inherent to LiDAR sensors.
Figure 26: Trajectory comparison generated by LIO-SAM, MA-LIO, and ORB-SLAM3, visualized using evo, alongside the ground truth trajectory from RTK.
Figure 27: Outdoor dataset trajectory in the x, y, and z coordinates.
TABLE V: Comparison of absolute pose error (APE) and relative pose error (RPE) across different SLAM algorithms.
LIO-SAM MA-LIO ORB3 mono imu ORB3 stereo imu
APE (m) 0.789 0.935 4.611 22.055
RPE (m) 1.307 0.114 0.036 0.041
X Future Work
This paper introduced the ShanghaiTech Mapping Robot and the outdoor dataset collected with it. The future work ahead of us can be split into two categories: further improving the robot and data collected as well as using the datasets for various exciting research areas.
We are currently working on upgrading the sensor setup of the robot by adding more event cameras (to stereo front and monocular left, right, back, up and down). We will also improve the data quality by filtering out the interfered points identified in Section VIII.
Then we will collect the definitive ShanghaiTech Robotic Dataset, with up to 2 hours indoor and outdoor sequences, augmented by data from other robot modalities (e.g. drones and quadrupeds). This dataset will also feature ground truth 3D point cloud data collected with a Faro 3D scanner. Data will also be collected at different times of the day, including during lunch rush-hour, with lots of students as dynamic objects. We also plan to collect datasets in other environments, such as industry, ancient architecture or agriculture settings.
In order to cover bigger outdoor areas we have been working on mounting the ShanghaiTech Mapping Robot on an electric flatbed car, increasing the speed of the data acquisition platform from 1 m/s to roughly 7 m/s [60]. Data from this setup will be included in the ShanghaiTech Robotic Dataset as well.
We may also pursue making the ShanghaiTech Mapping Robot fully autonomous in the future, to enable the capture of datasets along exactly the same trajectory, which might be useful for certain types of approaches [16], such as lifelong SLAM [61].
The datasets collected with our system will enable a wide range of research topics. The foremost application and one of the main motivations for building this robot is to evaluate SLAM algorithms in great depth. For this we have developed the SLAM Hive Benchmarking Suite [6, 7], a docker-based distributed system to run, evaluate and analyze SLAM systems with various datasets, based on their ground truth paths. The datasets collected with our robot open up the full potential of SLAM Hive, by providing high-resolution, high-frame rate data from various sensors in various configurations, thus, for the first time, enabling a thorough, automated comparison of SLAM algorithms of various flavours (e.g. LiDAR-based, RGB, RGB-D, event, radar, …) with various sensor configurations (e.g. front- vs. side-facing sensors), sensor combinations and algorithm configurations (e.g. number of particles, number of features, frame rate, resolution). We hope that this will give important insights into SLAM algorithms and provide a powerful tool for the design and evaluation of future SLAM approaches.
Down the line we furthermore plan to fully label the ground truth FARO point cloud. Together with the ground truth path of the robot we can then project those labels back to all sensor data from the robot. We hope that this can further drive the research in areas such as segmentation or object detection and tracking. The datasets can then also be used in an extended version of SLAM Hive to benchmark such kind of algorithms.
In general, given the data-richness of our dataset, we hope that researchers will use it to develop and test all kinds of interesting algorithms, especially w.r.t to the fusion of unconventional sensors, e.g. event with ultrasound and IMU data, or event with radar, etc.
Finally, we will also utilize the datasets to generate osmAG maps (open street map Area Graph) [62], which we have utilized for Robust Lifelong Indoor Localization [63] and to enable LLMs (Large Language Models) to interpret maps [64].
XI Conclusions
In this paper we have presented the ShanghaiTech Mapping Robot and the outdoor dataset collected with it. The robot features a never-seen before richness of high-quality sensors, including RGB and RGB-D cameras, IR and event cameras, LiDARs, IMUs, ultrasound and mmWave range sensors as well as an GNSS RTK receiver. A 16-node cluster was constructed to enable the collection of all of this data. The mechanical structure is very rigid to enable the precise extrinsic calibration of the sensors. A hardware-synchronization approach ensure the proper triggering and time-stamping of most of the sensor data. We have described the procedures of dataset collection and our post-processing steps. Our investigation of the interference of the active IR-based sensors showed that just two of the front LiDAR suffer from a limited amount of range readings with added noise.
TABLE VI: Sensor Data Statistics of Outdoor Dataset